# Signature-Informed Transformer for Asset Allocation

### v1

@@ 4.3 Module‑Level Contribution Experiments
To dissect the contribution of each architectural pillar of the Signature-informed Transformer (SIT), we conduct a comprehensive ablation study. For this analysis, each ablated variant is created by independently removing a single key component from the full SIT model, while all other hyperparameters are held constant. This module-drop protocol allows for a precise evaluation of each component’s marginal contribution. The variants evaluated are: (i) w/o CVaR, which replaces the Conditional Value-at-Risk objective with a risk-neutral objective of maximizing mean returns (ii) w/o Asset Attn, which disables the entire Signature-Informed Self-Attention mechanism across assets (iii) w/o Financial Bias, which removes the signature-derived bias term from the attention scores, reverting to a standard self-attention mechanism and (iv) w/o Gate $\gamma$, which removes the learnable gate $\gamma$ that scales the financial bias.
@@
w/o CVaR 0.5691 0.7057 0.3695 1.6409
w/o Asset Attn 0.5284 0.6576 0.3342 1.5381
w/o Financial Bias 0.6045 0.7590 0.3431 1.6801
w/o Gate $\gamma$ 0.5251 0.6489 0.3470 1.5470
@@
w/o CVaR 0.5923 0.7294 0.3606 1.6622
w/o Asset Attn 0.6268 0.7650 0.3298 1.6562
w/o Financial Bias 0.6047 0.7545 0.3224 1.6260
w/o Gate $\gamma$ 0.5831 0.7138 0.3305 1.5945
Table 2: Ablation study of SIT’s core components.
The results, summarized in Table 2, underscore the importance of each design choice. The most critical element is the decision-focused approach. When the Conditional Value-at-Risk (CVaR) loss is replaced with a standard risk-neutral objective (w/o CVaR), the Sharpe Ratio on the 40-asset universe falls from 0.6717 to 0.5691. This demonstrates that direct optimization for risk-adjusted outcomes is essential for producing stable allocations that are resilient to tail events.
The components of the Signature-Informed Self-Attention mechanism prove equally vital. Removing the asset-wise attention layer entirely (w/o Asset Attn) severely reduces the model’s ability to reason about portfolio structure, causing a steep performance drop (Sharpe of 0.5284). Furthermore, removing just the signature-based inductive bias (w/o Financial Bias), i.e. reverting to a standard attention mechanism, still leads to significant degradation (Sharpe of 0.6045). This confirms that injecting principled geometric knowledge of lead-lag structures (Theorem 3.1) is more effective than forcing the model to learn these relationships from scratch. Finally, removing the learnable gate $\gamma$ (w/o Gate $\gamma$) is highly detrimental (Sharpe of 0.5251), highlighting that the model must learn to dynamically modulate the influence of these financial priors (Theorem 3.2) to adapt to changing market regimes.
4.4 Sensitivity to Transaction Costs and Allocation Concentration $\tau$
We examine how proportional trading frictions and allocation concentration affect SIT. The concentration parameter $\tau>0$ is the softmax temperature in the allocation layer, $\mathbf{w}^{(k)}_{t}=\text{softmax}(\hat{\boldsymbol{\mu}}^{(k)}_{t}/\tau)$. Smaller $\tau$ concentrates capital, larger $\tau$ spreads it. We sweep $\tau\in\{0.8,...,1.4\}$. Transaction costs are one‑way proportional fees of $c\in{0,5,10}$ basis points (bps; $1$ bps $10^{-4}$) per dollar traded. All other settings follow the main evaluation: long‑only, fully invested, monthly ($k$-step) rebalancing on the 40‑ and 50‑asset universes, and a zero risk‑free rate for all risk‑adjusted metrics.
Figure 3 reports mean ($\pm$ std) Sharpe ratios for every $(\tau,c)$ pair, with the 40‑asset universe on the left and 50‑asset on the right. Two patterns are stable across universes. First, performance peaks at moderate dispersion, near $\tau\approx 1.3$. Second, frictions compress Sharpe roughly linearly over this range: moving from 0 to $10$ bps reduces Sharpe by about 0.030.030.040.04 at the optimum.
@@ 4.4 Sensitivity to Transaction Costs and Allocation Concentration $\tau$
Figure 3: Sharpe ratio sensitivity to transaction costs and allocation concentration ($\tau$). Values are mean ($\pm$ std). Left: 40 assets Right: 50 assets.
@@ 5 Conclusion
This work argues that effective quantitative portfolio management requires robust allocation policies, not just optimizing prediction accuracy. We introduce the Signature-informed Transformer (SIT), a novel framework using path signatures for rich feature representation, a signature-augmented attention mechanism for financial biases like lead-lag effects, and a training objective that directly minimizes portfolio Conditional Value-at-Risk. Our empirical results show that SIT decisively outperforms baselines, which often are harmed by objective mismatch and error amplification. SIT’s performance remains superior under realistic transaction costs, underscoring the importance of its calibrated, signature-based architecture. While tested on U.S. equity data, this framework could be extended to higher-frequency, global, multi-asset markets. Ultimately, SIT provides a blueprint for ML systems to progress from forecasting towards a more end-to-end, risk-aware capital allocation.
Appendix A Appendix. Notation
@@
Appendix B Appendix. Mathematical Proofs
Definition B.1.
@@
Theorem B.2.
(Strict Lead-Lag Implies Positive Second-Order Signature) Let $\mathbf{X}_{t}=(X^{1}_{t},X^{2}_{t})$ for $t\in[0,T]$ satisfy the strict lead-lag structure of Definition B.1. Then the second-level signature cross-term
@@
Let $\mathbf{X}_{t}=(X^{1}_{t},X^{2}_{t})_{t\in[0,T]}$ be a path of bounded variation with the strict lead-lag structure of Definition B.1. By this structure, there exists a partition $0=t_{0}<t_{1}<\cdots<t_{2N}=T$ such that on each interval $t_{2k-2},t_{2k-1}]$ only $X^{1}$ varies (while $X^{2}$ remains constant), and on the following interval $t_{2k-1},t_{2k}]$ only $X^{2}$ varies (while $X^{1}$ is constant). Moreover, at the synchronization times $t_{2k}$ both coordinates coincide, and no increment is zero.
Recall from Definition B.1 the common values at the synchronization points:
@@
Theorem B.3 (Positive directional derivative of attention weight).
@@
Appendix C Appendix. Implementation details
@@
Panel D. SIT (Ours)
@@
Appendix D Appendix. Why we choose CVaR?

### v2

@@ 3.3 Module‑Level Contribution Experiments
To dissect the contribution of each architectural pillar of the SIT, we conduct a comprehensive ablation study. For this analysis, each ablated variant is created by independently removing a single key component from the full SIT model, while all other hyperparameters are held constant. This module-drop protocol allows for a precise evaluation of each component’s marginal contribution. The variants evaluated are: (i) w/o CVaR, which replaces the Conditional Value-at-Risk objective with a risk-neutral objective of maximizing mean returns (ii) w/o Asset Attn, which disables the entire Signature-Informed Self-Attention mechanism across assets (iii) w/o Financial Bias, which removes the signature-derived bias term from the attention scores, reverting to a standard self-attention mechanism and (iv) w/o Gate $\gamma$, which removes the learnable gate $\gamma$ that scales the financial bias.
The results, summarized in Table 2, underscore the importance of each design choice. The most critical element is the decision-focused approach. When the Conditional Value-at-Risk (CVaR) loss is replaced with a standard risk-neutral objective (w/o CVaR), the Sharpe Ratio on the 40-asset universe falls from 0.6717 to 0.5691. This demonstrates that direct optimization for risk-adjusted outcomes is essential for producing stable allocations that are resilient to tail events. The components of the Signature-Informed Self-Attention mechanism prove equally vital. Removing the asset-wise attention layer entirely (w/o Asset Attn) severely reduces the model’s ability to reason about portfolio structure, causing a steep performance drop (Sharpe of 0.5284). Furthermore, removing just the signature-based inductive bias (w/o Financial Bias), i.e. reverting to a standard attention mechanism, still leads to significant degradation (Sharpe of 0.6045). This confirms that injecting principled geometric knowledge of lead-lag structures (Theorem 2.1) is more effective than forcing the model to learn these relationships from scratch. Finally, removing the learnable gate $\gamma$ (w/o Gate $\gamma$) is highly detrimental (Sharpe of 0.5251), highlighting that the model must learn to dynamically modulate the influence of these financial priors (Theorem 2.2) to adapt to changing market regimes.
@@
w/o CVaR 0.5691c 0.7057b 0.3695 1.6409b
w/o Asset Attn 0.5284c 0.6576c 0.3342b 1.5381c
w/o Financial Bias 0.6045b 0.7590b 0.3431c 1.6801c
w/o Gate $\gamma$ 0.5251 0.6489c 0.3470b 1.5470c
@@
w/o CVaR 0.5923c 0.7294c 0.3606b 1.6622c
w/o Asset Attn 0.6268c 0.7650b 0.3298 1.6562b
w/o Financial Bias 0.6047c 0.7545b 0.3224 1.6260b
w/o Gate $\gamma$ 0.5831c 0.7138c 0.3305 1.5945c
Table 2: Ablation study of SIT’s core components.Superscripts $b$ and $c$ indicate statistical significance at $p<0.05$ and $p<0.001$ in paired tests against SIT (Ours).
3.4 Are signatures effective at driving attention?
SIT perturbs asset‑axis attention logits by an additive, signature‑induced bias, $\text{Logits}_{k,h}=\frac{\mathbf{Q}_{k,h}\mathbf{K}_{k,h}^{\top}}{\sqrt{d_{k}}}+\gamma\mathbf{B}_{k,h}\in\mathbb{R}^{d\times d}$ where $\mathbf{B}_{k,h}$ aggregates the alignment equation (10) between the query’s informational need and the cross‑signature embedding. Theorem 2.2 predicts that increasing this alignment in the query direction strictly raises attention weight on the corresponding key when $\gamma>0$. Coupled with Theorem 2.1, which links persistent lead–lag to non zero second-order signatures, the method suggests a testable implication. Consequently, assets whose signatures are stronger should systematically attract more inbound attention. We formalize this implication by defining, at each decision time $t$, a per‑asset signature‑strength score $s_{t,j}=\frac{1}{HN_{H}d}\sum_{k=1}^{H}\sum_{h=1}^{N_{H}}\sum_{m=1}^{d}(\mathbf{B}_{k,h})_{m,j}$ and the corresponding inbound‑attention share $a_{t,j}=\frac{1}{HN_{H}d}\sum_{k=1}^{H}\sum_{h=1}^{N_{H}}\sum_{m=1}^{d}\alpha_{k,h,m\to j}$. We then compute the Pearson correlation between $\{s_{t,j}\}_{j=1}^{d}$ and ($\{a_{t,j}\}_{j=1}^{d}$) at each $t$ and examine the distribution of these correlations across the test horizon.
Figure 3: Distribution of correlations between signature strength and attention weights.
In Figure 3, the distribution is right‑skewed with means of 0.540 for the 40‑asset universe and 0.611 for the 50‑asset universe, indicating that stronger signature signals are associated with higher inbound attention mass. The heavy right tail shows frequent periods in which attention concentrates on assets whose signature‑derived relations are most salient, while the paucity of negative correlations rules out a degeneracy in which the bias is ignored by the attention mechanism. These empirical patterns directly instantiate the monotonicity predicted by Theorem 2.2. So, as alignment $b_{k,h,j,l}$ strengthens, the induced change in $\alpha_{j,l}$ is positive, and the learned $\gamma$ scales this effect without flipping its sign.
The financial significance of this correlation is twofold. First, $\mathbf{B}$ is not a static embedding. This couples dynamic queries $\mathbf{q}^{\text{dyn}}_{k,j}$ with cross‑signatures $\boldsymbol{\beta}_{i,j,l}$. Hence, the correlation becomes most apparent when the model routes information toward assets whose relational signatures are currently informative. Second, because SIT is trained solely through the portfolio‑level $\text{CVaR}_{\alpha}$ loss, the learned attention must improve tail‑aware allocations rather than just forecast error. The observed right‑skew therefore indicates that signatures are not merely present but actively utilized to amplify risk‑relevant dependencies. In Appendix  , Figure  7 overlays the learned gate $\gamma$ on portfolio drawdowns. We observe that higher values of $\gamma$ tend to cluster during volatile episodes. This suggests that SIT increases the weight of signature‑based priors precisely when cross‑asset relations are most informative and prediction noise is elevated. We therefore conjecture that this financial bias offers a plausible explanation for the results in Table 1. Specifically, it allows SIT to achieve the high risk-adjusted returns while maintaining a robust diversification profile comparable to the EWP.
3.5 Why PFL approach fail financial objectives?
Across all eight forecasting models, prediction alone does not translate into superior trading performance. As illustrated in Figure 4, which reports out‑of‑sample Sharpe ratios after CVaR optimization, the decision‑focused learning (blue) consistently dominates the prediction‑only approach (gray) across both the 40‑ and 50‑asset universes. Moreover, as observed in Figure 1, the gap widens in the 50‑asset universe, indicating that higher dimensionality amplifies the failure of the predict‑then‑optimize pipeline.
Figure 4: Out‑of‑sample Sharpe ratios after CVaR portfolio optimization on 40‑ and 50‑asset S&P100 universes.
@@ 3.5 Why PFL approach fail financial objectives?
Figure 5: Sharpe ratio sensitivity to transaction costs and allocation concentration ($\tau$). Left: 40 assets Right: 50 assets.
The failure of MSE-based approaches stems from the objective mismatch. See Appendix F for details on the two-stage implementation. Prediction‑focused training optimizes a surrogate that is misaligned with the financial goal. Therefore, prediction losses weight all errors equally and are blind to the downstream mapping from forecasts to actions. A model that is optimal for $L_{\text{pred}}$ need not be even approximately optimal for CVaR. Consequently, tiny cross‑sectional ranking errors induced by MSE training can be amplified by the optimizer, effectively flipping the identity of the largest weights. The evidence in Figure 4 and the mechanisms above explain why predict‑then‑optimize pipelines produce low and unstable Sharpe despite competitive $L_{\text{pred}}$. By differentiating through the portfolio layer and optimizing the risk metric of interest, decision‑focused learning reshapes the logits so that only forecast features that improve allocation under constraints and tails are amplified. This alignment both raises risk‑adjusted returns and tightens variability across runs, which we observe consistently across backbones and universes. We believe that aligning the training objective with the financial objective is necessary for turning predictive signals into reliable portfolios.
3.6 Sensitivity Analysis
We examine how proportional trading frictions and allocation concentration affect SIT. The concentration parameter $\tau$ is the softmax temperature in the allocation layer, $\mathbf{w}^{(k)}_{t}=\text{softmax}(\hat{\boldsymbol{\mu}}^{(k)}_{t}/\tau)$. Smaller $\tau$ concentrates capital, larger $\tau$ spreads it. We sweep $\tau\in\{0.8,...,1.4\}$. Transaction costs are one‑way proportional fees of $c\in{0,5,10}$ basis points per dollar traded. All other settings follow the main evaluation: long‑only, fully invested, monthly ($k$-step) rebalancing on the 40‑ and 50‑asset universes, and a zero risk‑free rate for all risk‑adjusted metrics. Figure 5 reports mean ($\pm$ std) Sharpe ratios for every $(\tau,c)$ pair, with the 40‑asset universe on the left and 50‑asset on the right. Two patterns are stable across universes. First, performance peaks at moderate dispersion, near $\tau\approx 1.3$. Second, frictions compress Sharpe roughly linearly over this range: moving from 0 to $10$ bps reduces Sharpe by about 0.030.030.040.04 at the optimum.
@@ 3.6 Sensitivity Analysis
Figure 6: Impact of temperature $\tau$ on portfolio diversification.
Also, as shown in Figure 6, increasing the softmax temperature $\tau$ monotonically reduces cross‑sectional concentration and increases diversification. Herfindahl–Hirschman Index $\mathrm{HHI}(\mathbf{w})=\sum_{j}w_{j}^{2}$ declines, while Shannon entropy $\mathcal{H}(\mathbf{w})=-\sum_{j}w_{j}\log w_{j}$ rises. For any fixed $\tau$, the 50‑asset universe achieves lower $\mathrm{HHI}$ and higher $\mathcal{H}$ than the 40‑asset universe, indicating a more diffuse allocation when the investable set is broader. These diagnostics provide an interpretable, one‑to‑one control of concentration through $\tau$, useful when desk policies cap the effective number of active lines or impose minimum diversification.
@@ 4 Conclusion
This work argues that effective quantitative portfolio management requires robust allocation policies, not just optimizing prediction accuracy. We introduce the Signature-informed Transformer, a novel framework using path signatures for rich feature representation, a signature-augmented attention mechanism for financial biases like lead-lag effects, and a training objective that directly minimizes portfolio Conditional Value-at-Risk. Our empirical results show that SIT decisively outperforms baselines, which often are harmed by objective mismatch and error amplification. SIT’s performance remains superior under realistic transaction costs, underscoring the importance of its calibrated, signature-based architecture. While tested on U.S. equity data, this framework could be extended to higher-frequency, global, multi-asset markets. Ultimately, SIT provides a blueprint for ML systems to progress from forecasting towards a more end-to-end, risk-aware capital allocation.
Appendix A Appendix. Related works
Deep Learning in Asset Allocation
The application of deep learning in quantitative trading has largely bifurcated into two distinct paradigms. The first, the classic Predict Focused Learning (PFL) pipeline, focuses on developing return-prediction models. In this stream of research, complex architectures map market data to future price movements. For instance, Transformers have been adapted to capture temporal dependencies in asset prices for return forecasting (Fischer and Krauss, 2018; Yoo et al., 2021; Lim et al., 2021). Some models employ Graph Neural Networks (GNNs) to explicitly model inter-asset relationships, such as sector correlations, to improve prediction accuracy (Xu et al., 2021; Duan et al., 2025). Despite their architectural novelty, these methods inherit the fundamental flaws of a decoupled approach (Lee et al., 2024b). They suffer from objective mismatch, as optimizing for prediction error (e.g., Mean Squared Error) does not guarantee profitable portfolio construction, and are susceptible to error amplification, where small prediction inaccuracies lead to drastically suboptimal and unstable allocations (Chung et al., 2022). A more promising direction, which we term Decision Focused Learning (DFL), seeks to overcome these limitations by training policies end-to-end. These models learn a direct mapping from market state to portfolio allocations, optimizing a true financial objective like a risk-adjusted return metric. Foundational work demonstrated how to embed financial operators, such as portfolio value and Sharpe ratio, within a deep network, making the entire strategy differentiable and trainable via gradient descent (Buehler et al., 2019; Zhang et al., 2020; Costa and Iyengar, 2023). Recent research has increasingly emphasized embedding practical portfolio constraints into the model training phase. Typical examples include prohibiting short selling, ensuring full investment (i.e., portfolio weights sum to one), and placing upper or lower bounds on individual asset allocations, all of which are incorporated directly into the model architecture or loss function (Lee et al., 2024a; Hwang et al., 2025a). While these end-to-end frameworks efficiently align the model’s training objective with financial goals, they often fall short in explicitly guiding the model to learn and utilize the diverse information present in multi-asset settings. This leaves a critical research gap. These models lack a strong financial inductive bias to explicitly represent the non-linear, path-dependent nature of price series and the geometric, time-local lead-lag relationships between assets. In our implementation, the predicted returns $\hat{\mu}$ serve only as internal logits for a differentiable allocation layer. All parameters are trained end-to-end solely through the portfolio-level CVaR objective, not a pointwise prediction loss, aligning with decision-focused learning. Our work addresses this gap by integrating the mathematical theory of path signatures directly into a transformer’s attention mechanism, creating an optimization-aware model that is architecturally designed to understand the underlying geometry of market dynamics. See (Lee et al., 2024b; Hwang et al., 2025b) for more detailed review of asset allocations
Transformer-Based Time Series Forecasting
The success of the Transformer architecture in natural language processing has inspired its widespread adoption for time series forecasting. The core innovation, the self-attention mechanism, allows these models to dynamically weigh the importance of all past time steps when predicting future values, enabling them to capture complex, long-range dependencies without the sequential processing limitations of recurrent neural networks(Vaswani et al., 2017; Li et al., 2019). To extend the receptive field without incurring the quadratic cost of full attention, a stream of variants introduce sparsity or hierarchical structure. For example, LogSparse(Li et al., 2019), ProbSparse(Zhou et al., 2021) and related kernels discard low-magnitude query–key interactions to achieve $\mathcal{O}(L\log L)$ complexity while retaining global context. From a more fundamental time series data perspective, Autoformer(Wu et al., 2021), FEDformer(Zhou et al., 2022) and ETSformer(Woo et al., 2022) decompose signals into trend–seasonality (or frequency-domain) components so that long-horizon patterns can be modeled additively and multiplicatively with reduced error accumulation. More recent PatchTST(Nie et al., 2022) and TimesNet(Wu et al., 2022) patch neighboring observations or convolve multi-scale windows before attention, embedding stronger inductive biases for periodicity and aliasing control. While these innovations alleviate the long-range dependency bottleneck, they remain largely data-agnostic. When applied to financial series they struggle with regime-dependent non-stationary, heavy-tailed noise, and asynchronous cross-asset lead–lag effects, causing attention scores to lock onto transient outliers and degrading out-of-sample robustness (Cartea et al., 2023; Cont, 2001; Miori and Cucuringu, 2022). Our approach departs from this paradigm by embedding each asset’s path in a Rough Path Signature space that is stable under time-reparameterization and robust to micro-structure noise, and by augmenting the attention logits with second-order cross-signature terms that encode the signed-area geometry underpinning lead–lag dynamics. Coupled with scenario-based optimization to hedge against structural breaks, SIT addresses both the generic long-range dependency problem and the finance-specific pathologies that limit existing Transformer forecasters.
Path Signatures in Time Series and Finance
The path signature, originating from Rough Path Theory, offers a non-parametric and faithful representation of streamed data by summarizing the geometry of a path as a sequence of iterated integrals (Lyons, 1998). A key property is its universality: any continuous function on the space of paths can be arbitrarily well-approximated by a linear function of the signature’s terms, making it a powerful basis for feature extraction (Chevyrev and Kormilitzin, 2016). In practice, the signature is truncated at a finite order $M$, yielding a vector $\mathrm{Sig}^{M}(\mathbf{X})$ that is robust to irregular sampling due to its invariance to time reparameterization. However, this truncation introduces a trade-off, as the feature dimension grows exponentially with the order $M$ and polynomially with the path dimension $d$, posing a significant computational burden. This challenge has motivated alternatives like signature kernels, which compute inner products in the high-dimensional feature space implicitly, avoiding explicit feature construction (Király and Oberhauser, 2019). In machine learning, signatures provide a potent inductive bias for modeling systems with path-dependent memory. The most direct application involves using truncated signatures as static input features for standard models (Gyurkó et al., 2013). More sophisticated integrations are found in continuous-time models like Neural Controlled Differential Equations (CDEs), which learn a vector field that is controlled by the input path, effectively modeling the system’s response to a driving signal (Kidger et al., 2020). For finance, a crucial insight arises from the signature’s geometry: the second-order terms of a joint signature over two asset paths precisely encode their signed area, a direct and robust measure of their temporal lead-lag relationship (Lyons and McLeod, 2022). This property has been successfully leveraged to build kernels for detecting asymmetric dependencies between financial instruments, offering a principled alternative to traditional correlation measures (Bonnier et al., 2019). Recent advancements extend this to attention mechanisms. the Rough Transformer (Moreno-Pino et al., 2024) introduces multi-view signature attention to operate directly on continuous-time representations. Also, applications to finance span volatility/return modeling, derivatives, and market microstructure. Early studies extracted signature coordinates to forecast realized volatility and to detect temporal asymmetries (Gyurkó et al., 2013). In options, signatures parameterize no‑arbitrage dynamics and enable data‑driven pricing/hedging (Arribas et al., 2020), including transformer‑style encoders fed with log/signatures (Tong et al., 2023). A crucial geometric motif is the second‑order signed area,
$A(X^{i},X^{j})=\int X^{i}\,dX^{j}-\int X^{j}\,dX^{i},$ (21)
which encodes temporal asymmetry and lead–lag; signature kernels exploit this to compare pairs or small baskets of assets (Chevyrev and Kormilitzin, 2016; Király and Oberhauser, 2019). Our architecture operationalization this motif at scale: SIT injects cross‑asset signature information as a dynamic, query‑conditioned bias inside attention, so that pairwise signed‑area evidence modulates which assets attend to which others at each decision point (cf. Theorem 2.1). While signatures mitigate non‑stationarity and encode higher‑order interactions, they incur truncation bias and can suffer from a curse of dimensionality as either degree $M$ or the number of assets grows; kernelization trades feature savings for quadratic kernel costs (Salvi et al., 2021; Bonnier et al., 2019). Compared with state‑space or transformer baselines, signatures offer complementary bias—geometric invariances and lead–lag structure—rather than longer receptive fields alone. Prior signature‑based works typically (i) use signatures as fixed inputs or kernels outside the attention mechanism and (ii) optimize predictive losses, not portfolio objectives (Gyurkó et al., 2013; Tong et al., 2023; Bonnier et al., 2019). SIT differs by coupling signature‑augmented, cross‑asset attention with end‑to‑end CVaR optimization for long‑only, fully‑invested portfolios, aligning representation, interaction, and objective (Buehler et al., 2019).
Appendix B Appendix. Notation
@@
Appendix C Appendix. Mathematical Proofs
Definition C.1.
@@
Theorem C.2.
(Strict Lead-Lag Implies Positive Second-Order Signature) Let $\mathbf{X}_{t}=(X^{1}_{t},X^{2}_{t})$ for $t\in[0,T]$ satisfy the strict lead-lag structure of Definition C.1. Then the second-level signature cross-term
@@
Let $\mathbf{X}_{t}=(X^{1}_{t},X^{2}_{t})_{t\in[0,T]}$ be a path of bounded variation with the strict lead-lag structure of Definition C.1. By this structure, there exists a partition $0=t_{0}<t_{1}<\cdots<t_{2N}=T$ such that on each interval $t_{2k-2},t_{2k-1}]$ only $X^{1}$ varies (while $X^{2}$ remains constant), and on the following interval $t_{2k-1},t_{2k}]$ only $X^{2}$ varies (while $X^{1}$ is constant). Moreover, at the synchronization times $t_{2k}$ both coordinates coincide, and no increment is zero.
Recall from Definition C.1 the common values at the synchronization points:
@@
Theorem C.3 (Positive directional derivative of attention weight).
@@
Appendix D Appendix. Implementation details
@@
Panel D. RFormer
Parameter Values
Embedding_Dim 8, 16, 32
E_LAYERS_LIST 1, 2
N_HEADS_LIST 2, 4, 8
Sig_Level 2, 3
Panel E. SIT (Ours)
@@
Appendix E Appendix. Why we choose CVaR?